OpenStudy (anonymous):

Find all polar coordinates of point P where P = (9, - pi/5)

OpenStudy (dan815):

is that given in cartesian or polar form that point

OpenStudy (dan815):

if that is already polar then, u can say 9,-pi/5 +k*2pi, where k are all the integers

OpenStudy (dan815):

where k is

OpenStudy (dan815):

of its caretsian point use the fact that the x^2+y^2=r^2 to find r and then angle = arctan(y/x)

OpenStudy (dan815):

and then ya.. carry on.. see how far u get with that first

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

sorry the answers to the question are a.)(9, -pi/ 5 + 2nπ) or (-9, -pi/ 5 + 2nπ) b.)(9, -pi/5 alt tag: negative pi divided by 5 + (2n + 1)π) or (-9,-pi/5 + 2nπ) c.)(9, -pi/ 5 + 2nπ) or (-9, -pi/ 5 + (2n + 1)π) and d.) (9, -pi/ 5 + 2nπ) or (9, pi/y 5 + (2n + 1)π)

OpenStudy (dan815):

ok

OpenStudy (dan815):

so what they gave you is polar form already

OpenStudy (dan815):

withevery revolution u will be back at the same point

OpenStudy (dan815):

revolution is a multiple of 2pi

OpenStudy (anonymous):

okay i was thinking a but I don't know

OpenStudy (dan815):

there is a y in d is that suppose to be there

OpenStudy (anonymous):

no